Orwell building supplies' last dividend was $1.75. Its dividend growth rate is expected to be constant at 10.00% for 2 years, after which dividends are expected to grow at a rate of 6% forever. Its required return (rs) is 12%. What is the best estimate of the current stock price? Select the correct answer. a. $27.83 b. $33.23 c. $31.88 d. $30.53 e. $29.18
Answer #1

D1=(1.75*1.1)=1.925

D2=(1.925*1.1)=2.1175

Value after year 2=(D2*Growth rate)/(Required return-Growth rate)

=(2.1175*1.06)/(0.12-0.06)

=37.4091667

Hence current price=Future dividend and value*Present value of discounting factor(rate%,time period)

=1.925/1.12+2.1175/1.12^2+37.4091667/1.12^2

=$33.23(Approx).